OpenStudy (anonymous):

What is the derivative of 2x^3

OpenStudy (anonymous):

by hand or power rule?

OpenStudy (anonymous):

by hand

OpenStudy (anonymous):

please

OpenStudy (anonymous):

2x^3 rule: nx^n-1 n=3 so (3)2x^3-1 =6x^2

OpenStudy (anonymous):

Sorry I need it by hand suka

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{2(x+h)^3-2x^3}{h}\] \[=2\lim_{h\to 0}\frac{(x+h)^3-x^3}{h}\] \[=2\lim_{h\to 0}\frac{x^3+3x^2h+3xh^2+h^3-x^3}{h}\] get rid of the \(x^3\) up top, factor out an \(h\) to cancel with the \(h\) in the denominator, then replace \(h\) by zero

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

yw, hope steps are clear, and don't forget that 2 out front

OpenStudy (anonymous):

is it 3*2x^2=6x^2